OpenStudy (anonymous):

Given A = 67o, B = 13o, and b = 20.3, which of the following can be used to solve the triangle? a. Law of Sines b. Law of Cosines c. Hero's Formula d. Cannot be solved.

OpenStudy (anonymous):

those are degrees

OpenStudy (anonymous):

cosine formula is a good option

OpenStudy (anonymous):

Ah. Law of sines: \[\frac{\sin(a)}{A}=\frac{\sin(b)}{b}\] Law of cosines is \[c^2 = a^2+b^2 - 2ab\cos(C)\] So what are you given that can be used here?

OpenStudy (anonymous):

cannot be solved

OpenStudy (anonymous):

Sorry the law of sines should have been: \[\frac{\sin(A)}{a}=\frac{\sin(B)}{b}\]

OpenStudy (anonymous):

the angles

OpenStudy (anonymous):

oh and side b

OpenStudy (anonymous):

Yes it can. You have 1 side and three angles

OpenStudy (anonymous):

so law of sines?

OpenStudy (anonymous):

And i flipped all my capitol letters with the convention. Yes, Multiple uses of Law of sines would solve the triangle. You would have to do a 180-a-b to get angle c though...